Tohban Report 2015-07-01
From RHESSI Wiki
|Start Date:||24 Jun 2015|
|End Date:||01 Jul 2015|
|List all reports|
Solar activity was rather low this week. Currently there are three active regions that just showed up on the east limb. Activity is not expected to be so high next couple of days. The major flare this week was a M7.2 that RHESSI observed during its decay phase.
How many GOES flares occurred?
Flares above B, C, M, X class were 19 19 1 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 10 12 1 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 301 / 39 over the time range 23-Jun-15 30-Jun-15
SSR fill was pretty low this week, usually below 40%. On June 29 the were some spikes in the SSR filling that I would like to discuss.
Normal as far as the Tohban can tell.
En general detectors were stable this week, no changes were needed. Light leak is still observed on the Front segment of detector 2. D8 and D9 front livetime are still showing low values.
Last Wednesday June 24 we increased the threshold used to put the thin attenuator IN (A0 to A1). We went from a value of 5.1 % to 10 % That change took effect 2015-175-22:12:56 UTC. We agreed in to perform this experiment for about one month, it means that by July 22 we should asses how well this changed worked.
RHESSI has spun up between 2015-180-21:49:44 and 2015-180-23:41:52 UTC. We spun up from 14.41 to 14.99 RPM.
We're still in Active/Vigorous decimation, and still taking events at night.
|Night time data (fronts)||Taking night events|
|Night time data (rears)||Taking night events|
|Require extra passes?||No|
|Attenuator operation||Performing an experiment|